OpenStudy (anonymous):

Which point is on the graph of y = –2cos x? A. (2, 0) B. (0, -2) C. (2, -2) D. (0,0)

OpenStudy (anonymous):

@whpalmer4 or @jim_thompson5910 can either of you help with this?

jimthompson5910 (jim_thompson5910):

check each answer choice

jimthompson5910 (jim_thompson5910):

so let's check choice A

jimthompson5910 (jim_thompson5910):

y = -2*cos(x) 0 = -2*cos(2) ... plug in x = 2 and y = 0 0 = -2*0.9993908270191 0 = -1.9987816540382 that equation is false, so (2,0) is NOT a point on the graph of y = -2*cos(x)

jimthompson5910 (jim_thompson5910):

you do the same for the other choices until you get a true equation

OpenStudy (anonymous):

I got B

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

B is correct